# StiffGIPC: Advancing GPU IPC for Stiff Affine-Deformable Simulation

### v3

@@ Abstract.
Incremental Potential Contact (IPC) is a widely used, robust, and accurate method for simulating complex frictional contact behaviors. However, achieving high efficiency remains a major challenge, particularly as material stiffness increases, which leads to slower Preconditioned Conjugate Gradient (PCG) convergence, even with the state-of-the-art preconditioners. In this paper, we propose a fully GPU-optimized IPC simulation framework capable of handling materials across a wide range of stiffnesses, delivering consistent high performance and scalability with up to 10$\times$ speedup over state-of-the-art GPU IPC methods. Our framework introduces three key innovations: 1) A novel connectivity-enhanced Multilevel Additive Schwarz (MAS) preconditioner on the GPU, designed to efficiently capture both stiff and soft elastodynamics and improve PCG convergence at a reduced preconditioning cost. 2) A $C^{2}$-continuous cubic energy with an analytic eigensystem for inexact strain limiting, enabling more parallel-friendly simulations of stiff membranes, such as cloth, without membrane locking. 3) For extremely stiff behaviors where elastic waves are barely visible, we employ affine body dynamics (ABD) with a hash-based two-level reduction strategy for fast Hessian assembly and efficient affine-deformable coupling. We conduct extensive performance analyses and benchmark studies to compare our framework against state-of-the-art methods and alternative design choices. Our system consistently delivers the fastest performance across soft, stiff, and hybrid simulation scenarios, even in cases with high resolution, large deformations, and high-speed impacts. Our framework will be fully open-sourced upon acceptance.
@@ 1. Introduction
In summary, there is still plenty of room for optimizing linear solver preconditioners, strain limiting, global Hessian matrix assembly, etc., for realizing a highly GPU-optimized IPC framework that can efficiently simulate large-scale affine-deformable coupled scenarios. In this paper, we propose such a framework, achieving up to 10×10$\times$10 × speedup compared to GIPC via the following 4 major innovations:
@@
In section 7, we perform extensive and rigorous performance analyses and benchmark studies to validate our framework and compare it to state-of-the-art GPU IPC systems and alternative design choices that may seem reasonable but suffer from suboptimal performance in practice. Our framework exhibits the fastest performance in soft, stiff, and hybrid simulation scenarios, even with high resolution, extreme deformation, and high-speed impacts. Our system will be fully open-sourced upon acceptance.
@@ 5.1. Formulation
We follow Kim (2020) and use $\sqrt{I_{5}(\mathbf{F},\mathbf{n})}$ to define our inexact strain-limiting energy, as it effectively represents the stretch factor in the direction $\mathbf{n}$. Directly incorporating $I_{5}$ into the barrier-based strain-limiting model of Li et al. (2021) enables reusing the eigenanalysis from Kim (2020). However, backtracking-based line search filtering is still required since no analytic bound for feasible step sizes is available for the strain. This method is inefficient for parallelization because its iterative halving process, applied to each triangle, requires the global linear search iterations to continue until all triangles satisfy the strain limit. Additionally, because the barrier function diverges rapidly near the strain limit, solver performance becomes highly sensitive to the barrier stiffness, complicating parameter setting. To address this, we apply a cubic penalty function to inexactly handle the inequality constraint imposed by the strain limit. This approach allows for slight constraint violations while providing stretching resistance on the same scale as practical settings.
@@
(10) $\Psi_{SL}^{u}=\lambda^{\prime}a_{t}S(I_{5}(\mathbf{F},\mathbf{n}_{u}))(\sqrt{I_{5}(\mathbf{F},\mathbf{n}_{u})}-1)^{3},$
where $\lambda^{\prime}$ is the stiffness and $S$ is a piecewise function defined as
(11) $S(s)=\begin{cases}1,&s>1,\\ 0,&0\leq s\leq 1,\end{cases}$
and so the inexact strain-limiting force is only applied when the cloth is stretched. This idea aligns with the tension field theory (Steigmann, 1990; Skouras et al., 2014; Chen et al., 2023), which deactivates stiffness under compression to obtain a convex energy. Additionally, Kikuuwe et al. (2009) activates a cubic term under compression to mitigate the element inversion issue.
Equivalently, Equation 10 can also be written as
$\Psi_{SL}^{u}=\lambda^{\prime}a_{t}\max(\sqrt{I_{5}(\mathbf{F},\mathbf{n}_{u})}-1,0)^{3}.$
@@
This energy is well-defined even if the strain limit is exceeded, eliminating the need for line search filtering. Given that the Young’s modulus of real fabrics ranges from $1MPa$ to $10MPa$ (Penava et al., 2014) when the cloth is stretched, we set our $\lambda^{\prime}=5MPa$ by default, as it provides stretching resistance at the same scale when the cloth approaches the strain limit, e.g., $5\%$. Thus, our model prevents over-elongation in practical cloth manipulation scenarios. For $\Psi_{stretch}$, we can now safely use a small stiffness ($\lambda=0.05MPa$ in all our examples) to avoid membrane locking. Refer to Figure 8 for a plot of our composed membrane energy in the $\mathbf{n}_{u}$ direction. Since the stiffness for the shearing term $\Psi_{shear}$ primarily depends on fabric type and, based on our observations, has minimal impact on the over-elongation or membrane locking issues, we simply set it to 30% of the stretching stiffness and apply the original eigenanalysis from Kim (2020) to compute its PSD Hessian.
@@ 5.2. Eigenanalysis
Thus, we can directly compute the PSD Hessian matrix of $\Psi_{SL}$ in the $\mathbf{n}_{u}$ and $\mathbf{n}_{v}$ directions separately, applying the chain rule with $I_{5}$ as the intermediate variable:
$\mathbf{H}_{SL,*}$ $=3S(I_{5}(\mathbf{F},\mathbf{n}_{*}))\left(1-\frac{1}{\sqrt{I_{5}(\mathbf{F},\mathbf{n}_{*})}}\right)(\sqrt{I_{5}(\mathbf{F},\mathbf{n}_{*})}-1)\mathbf{H}_{*}$
(17) $\,\,+\frac{3S(I_{5}(\mathbf{F},\mathbf{n}_{*}))(I_{5}(\mathbf{F},\mathbf{n}_{*})-1)}{I_{5}(\mathbf{F},\mathbf{n}_{*})^{3/2}}(\mathbf{f}_{*}\mathbf{f}_{*}^{T}).$
@@ 6.2. Hash-based Parallel Reduction
The first two primitives are standard in parallel computing. For warp-level segmented reduction, we use the CUB [^2] https://nvidia.github.io/cccl/cub/ intrinsic function HeadSegmentedReduce(values, tags). Consider a simple example with a warp size of 8: if $\emph{{values}}=[1,1,1,1,1,1,1,1]$ and $\emph{{tags}}=[1,0,0,1,0,0,1,0]$, the output will be $\emph{{outPut}}=[3,\bullet,\bullet,3,\bullet,\bullet,2,\bullet]$, where $\bullet$ indicates invalid entries. Here, the tags array specifies start and end positions for reduction in values, with 1 marking the start. Reduction results (valid entries in outPut) are then written to the result array. The overall procedure, FastHashReduction, is detailed in Algorithm 2. First, the map array is constructed in parallel, mapping duplicate keys to their unique key indices to ensure reduction results are written to the correct locations (lines 1–8). The FastSegmentReduction algorithm (Algorithm 3) then efficiently computes the matrix.
@@
5:     $R$triangleright an array of reduced values
@@
In Algorithm 2, we create a mapping array $O$ to associate each key-value pair with its target memory location, as illustrated in Figure 10. Array $O$ is also used to generate tag information for the CUB intrinsic function HeadSegmentedReduce(values, tags). The tags are determined by comparing adjacent entries in $O$, as described in Algorithm 3. Here, $b^{-}$, $b$, and $b^{+}$ represent the previous, current, and next unique key indices, respectively. The IsValid flag checks whether the GlobalThreadId falls within valid data range. If $b^{-}\neq b$, it marks the start of a new segment. The IsCrossWarp flag indicates if the segment spans multiple warps, necessitating atomic operations to avoid data corruption; otherwise, direct writing is safe. These flags are set per thread, but only the leading thread (head thread) can write to the output array. Therefore, we merge flags across threads in each segment: a head thread is valid if any thread in the segment is valid and marked as cross-warp if any spans multiple warps. This is achieved via segmented warp reduction, using an OR operation for merging (see lines 5-28).
@@
2:     $O$triangleright The result in line 8 of 2
@@ 6.3. Symmetric Reduce-By-Key (SRBK) SpMV
Our SRBK SpMV and FastSegmentReduction algorithms have a similar structure, both using a ”Reduction By Key” parallel primitive. Sorted row indices serve as hash keys, while the product of the block matrix and the input vector represents the value to be reduced (lines 5-9). The main difference is in how we access the lower triangular part of the matrix, where we atomically add the multiplication results to the output vector (line 11). The data from the lower triangular part is directly transposed from the upper part, reducing global memory access by nearly half, and thus significantly enhancing performance.
@@ 7.1. Ablation and Comparative Studies
109K, 80K, 200K CIPC (CEMAS+SRBK) 104 202 21 91 6 424 2.89e5 26.25 69 -
@@
105K, 10K, 202K CIPC (CEMAS+SRBK) 219 812 59 329 10 1429 1.75e6 38.76 180 -

### v4

@@ Abstract.
Incremental Potential Contact (IPC) is a widely used, robust, and accurate method for simulating complex frictional contact behaviors. However, achieving high efficiency remains a major challenge, particularly as material stiffness increases, which leads to slower Preconditioned Conjugate Gradient (PCG) convergence, even with the state-of-the-art preconditioners. In this paper, we propose a fully GPU-optimized IPC simulation framework capable of handling materials across a wide range of stiffnesses, delivering consistent high performance and scalability with up to 10$\times$ speedup over state-of-the-art GPU IPC methods. Our framework introduces three key innovations: 1) A novel connectivity-enhanced Multilevel Additive Schwarz (MAS) preconditioner on the GPU, designed to efficiently capture both stiff and soft elastodynamics and improve PCG convergence at a reduced preconditioning cost. 2) A $C^{2}$-continuous cubic energy with an analytic eigensystem for inexact strain limiting, enabling more parallel-friendly simulations of stiff membranes, such as cloth, without membrane locking. 3) For extremely stiff behaviors where elastic waves are barely visible, we employ affine body dynamics (ABD) with a hash-based two-level reduction strategy for fast Hessian assembly and efficient affine-deformable coupling. We conduct extensive performance analyses and benchmark studies to compare our framework against state-of-the-art methods and alternative design choices. Our system consistently delivers the fastest performance across soft, stiff, and hybrid simulation scenarios, even in cases with high resolution, large deformations, and high-speed impacts.
@@ 1. Introduction
In summary, there is still plenty of room for optimizing linear solver preconditioners, strain limiting, global Hessian matrix assembly, etc., for realizing a highly GPU-optimized IPC framework that can efficiently simulate large-scale affine-deformable coupled scenarios. In this paper, we propose such a framework, achieving up to 10×10$\times$10 × speedup compared to GIPC via the following 3 major innovations:
@@
In section 7, we perform extensive and rigorous performance analyses and benchmark studies to validate our framework and compare it to state-of-the-art GPU IPC systems and alternative design choices that may seem reasonable but suffer from suboptimal performance in practice. Our framework exhibits the fastest performance in soft, stiff, and hybrid simulation scenarios, even with high resolution, extreme deformation, and high-speed impacts.
@@ 5.1. Formulation
We follow Kim (2020) and use $\sqrt{I_{5}(\mathbf{F},\mathbf{n})}$ to define our inexact strain-limiting energy, as it effectively represents the stretch factor in the direction $\mathbf{n}$. Directly incorporating $I_{5}$ into the barrier-based strain-limiting model of Li et al. (2021) enables reusing the eigenanalysis from Kim (2020). However, backtracking-based line search filtering is still required since no analytic bound for feasible step sizes is available for the strain. This method is inefficient for parallelization because its iterative halving process, applied to each triangle, requires the global line search iterations to continue until all triangles satisfy the strain limit. Additionally, because the barrier function diverges rapidly near the strain limit, solver performance becomes highly sensitive to the barrier stiffness, complicating parameter setting. To address this, we apply a cubic penalty function to inexactly handle the inequality constraint imposed by the strain limit. This approach allows for slight constraint violations while providing stretching resistance on the same scale as practical settings.
@@
(10) $\Psi_{SL}^{u}=\lambda^{\prime}a_{t}\max(\sqrt{I_{5}(\mathbf{F},\mathbf{n}_{u})}-1,0)^{3},$
where $\lambda^{\prime}$ is the stiffness, and the $\max$ operator ensures that the inexact strain-limiting force is only applied when the cloth is stretched. This idea aligns with the tension field theory (Steigmann, 1990; Skouras et al., 2014; Chen et al., 2023), which deactivates stiffness under compression to obtain a convex energy. Additionally, Kikuuwe et al. (2009) activates a cubic term under compression to mitigate the element inversion issue.
@@
This energy is well-defined even if the strain limit is exceeded, eliminating the need for line search filtering. Given that the Young’s modulus of real fabrics ranges from $1MPa$ to $10MPa$ (Penava et al., 2014) when the cloth is stretched, we set our $\lambda^{\prime}=5MPa$ by default, as it provides stretching resistance at the same scale when the cloth approaches the strain limit, e.g., $5\%$. Thus, our model prevents over-elongation in practical cloth manipulation scenarios.
For $\Psi_{stretch}$, we can now safely use a small stiffness ($\lambda=0.05MPa$ in all our examples) to avoid membrane locking. Refer to Figure 8 for a plot of our composed membrane energy in the $\mathbf{n}_{u}$ direction. Since the stiffness for the shearing term $\Psi_{shear}$ primarily depends on fabric type and, based on our observations, has minimal impact on the over-elongation or membrane locking issues, we simply set it to 30% of the stretching stiffness and apply the original eigenanalysis from Kim (2020) to compute its PSD Hessian.
@@ 5.2. Eigenanalysis
Thus, when the the cloth is stretched, we can directly compute the PSD Hessian matrix of $\Psi_{SL}$ in the $\mathbf{n}_{u}$ and $\mathbf{n}_{v}$ directions separately, applying the chain rule with $I_{5}$ as the intermediate variable:
$\mathbf{H}_{SL,*}$ $=3\left(1-\frac{1}{\sqrt{I_{5}(\mathbf{F},\mathbf{n}_{*})}}\right)(\sqrt{I_{5}(\mathbf{F},\mathbf{n}_{*})}-1)\mathbf{H}_{*}$
(16) $\,\,+\frac{3(I_{5}(\mathbf{F},\mathbf{n}_{*})-1)}{I_{5}(\mathbf{F},\mathbf{n}_{*})^{3/2}}(\mathbf{f}_{*}\mathbf{f}_{*}^{T}).$
@@ 6.2. Hash-based Parallel Reduction
The first two primitives are standard in parallel computing. For warp-level segmented reduction, we use the CUB [^2] https://nvidia.github.io/cccl/cub/ intrinsic function HeadSegmentedReduce(values, tags). Consider a simple example with a warp size of 8: if $\emph{{values}}=[1,1,1,1,1,1,1,1]$ and $\emph{{tags}}=[1,0,0,1,0,0,1,0]$, the output will be $\emph{{outPut}}=[3,\bullet,\bullet,3,\bullet,\bullet,2,\bullet]$, where $\bullet$ indicates invalid entries. Here, the tags array specifies start and end positions for reduction in values, with 1 marking the start. Reduction results (valid entries in outPut) are then written to the result array. The overall procedure, FastHashReduction, is detailed in Algorithm 2. First, the map array $O$ is constructed in parallel, mapping duplicate keys to their unique key indices to ensure reduction results are written to the correct locations (lines 2–6). The FastSegmentReduction algorithm (Algorithm 3) then efficiently computes the matrix.
@@
5:     $\mathbf{AG}$triangleright global Hessian matrix
@@
In Algorithm 2, we create a mapping array $O$ to associate each key-value pair with its target memory location, as illustrated in Figure 10. Array $O$ is also used to generate tag information for the CUB intrinsic function HeadSegmentedReduce(values, tags). The tags are determined by comparing adjacent entries in $O$, as described in Algorithm 3. Here, $b^{-}$ and $b$ represent the previous and current unique key indices, respectively. A new segment begins when $b^{-}\neq b$ or when the corresponding thread is the head of its warp. This is indicated using the IsHead flag. This flag is set per thread, but only the thread with the flag equal to $1$ is allowed to write to the output array.
@@
2:     $O$triangleright The result in line 6 of 2
@@ 6.3. Symmetric Reduce-By-Key (SRBK) SpMV
Our SRBK SpMV and FastSegmentReduction algorithms have a similar structure, both using a ”Reduction By Key” parallel primitive. Sorted row indices serve as hash keys, while the product of the block matrix and the input vector represents the value to be reduced (lines 4-23). The main difference is in how we access the lower triangular part of the matrix, where we atomically add the multiplication results to the output vector (line 14). The data from the lower triangular part is directly transposed from the upper part, reducing global memory access by nearly half, and thus significantly enhancing performance.
@@ 7.1. Ablation and Comparative Studies
109K, 80K, 200K CIPC (CEMAS+SRBK) 104 202 21 91 6 424 2.89e5 26.25 69 N/A
@@
105K, 10K, 202K CIPC (CEMAS+SRBK) 219 812 59 329 10 1429 1.75e6 38.76 180 N/A
@@ 8. Conclusion and Discussion
Acknowledgements.
This work was supported in part by the Junior Faculty Startup Fund of Carnegie Mellon University. It was also partially funded by the Research Grants Council of Hong Kong (Ref: 17210222), and by the Innovation and Technology Commission of the HKSAR Government under the ITSP-Platform grant (Ref: ITS/335/23FP) and the InnoHK initiative (TransGP project). Part of the research was conducted in the JC STEM Lab of Robotics for Soft Materials, funded by The Hong Kong Jockey Club Charities Trust. We thank the reviewers for their detailed and insightful feedback, and we are especially grateful to Xinyu Lu for his important contributions to the system implementation.